Simplify the expression. 1) 9x-2-2x+ 5 A) 10 в) 10x C) 11x +3 D) 7x +3 2) 7x -4x -x 12 A) 3x - 12 B) 3x -x-12 C) 2x -12 D) 3x -13 3)-12 -(5y-10) A) -5y +2 B) -5y-2 C) -5y+ 22 D) -5y-22 4) (8x 13) (3x -1) A) 5x 12 4) B) 6x 13 C) 6x- 12 D) 5x 13 5) 5) -(10v -6)+ 10(2v+ 10) A) 10v+ 106 B) -10v+ 94 C) 30v + 16 D) 30v+4 6) 6) 8(3x2-2) -6(x2 -1) A) -24x2-10 B) 18x2- 10 C) 18x2- 22 D) 1sx2-3 7)

Please help me with numbers 1-5 thanks!
Answer #1

Answer :

1) Consider the expression

9x - 2 - 2x + 5 = ( 9 - 2 )x - 2 + 5

= 7x + 3

2) Consider the expression

7x - 4x -x - 12 = (7 - 4 - 1)x - 12

= 2x - 12

3) Consider the expression

- 12 - ( 5y - 10 ) = - 12 - 5y + 10

= - 5y - 2

4) Consider the expression

(8x - 13) - (3x - 1) = (8 - 3)x - 13 + 1

= 5x - 12

5) Consider the expression

- (10v - 6) + 10(2v + 10) = -10v + 6 + 20v + 100

= 10v + 106

6) Consider the expression

8(3x2 - 2) - 6(x2 - 1) = 24x2 - 16 - 6x2 + 6

= (24 - 6)x2 - 16 + 6

= 18x2 - 10
Answer #2
1) 9x - 2 - 2x + 5 = ( 9 - 2 )x - 2 + 5 = 7x + 3 2) 7x - 4x -x - 12 = (7 - 4 - 1)x - 12 = 2x - 12 3) - 12 - ( 5y - 10 ) = - 12 - 5y + 10 = - 5y - 2 4) (8x - 13) - (3x - 1) = (8 - 3)x - 13 + 1 = 5x - 12 5) - (10v - 6) + 10(2v + 10) = -10v + 6 + 20v + 100 = 10v+106
answered by: Pabasara